# A Case Study of Cross-Lingual Zero-Shot Generalization for Classical Languages in LLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) (Brown, 2020; Ouyang et al., 2022; Touvron et al., 2023) are known to generalize across various tasks using data from languages present in their pre-training phase, even when not present in instruction tuning (Wang et al., 2022; Muennighoff et al., 2023; Han et al., 2024). Previous work has demonstrated generalization to several low-resource languages via few-shot in-context learning (Cahyawijaya et al., 2024). In this study, we focus on zero-shot generalization to natural language understanding (NLU) tasks for three Indo-European classical languages—Sanskrit (san), Ancient Greek (grc), and Latin (lat)—with a primary focus on Sanskrit. These languages represent a unique category of low-resource languages: despite scarce data for downstream NLU tasks, they have rich ancient literature available in digitized formats (Goyal et al., 2012; Berti, 2019) and exert significant influence on the vocabulary and narrative structures of better-resourced languages (e.g., Latin contributes approximately 28% of English vocabulary (Finkenstaedt and Wolff, 1973)). Moreover, the high inflection these languages present a challenge for evaluating LLM generalization.
To investigate these issues, we conduct two sets of zero-shot experiments using gpt-4o (OpenAI, 2024, 2023), llama-3.1-405b-instruct (Dubey et al., 2024), and their smaller variants. First, we assess performance on two NLU tasks with available datasets for all three languages, namely, named entity recognition (NER) and machine translation to English (MT). We observe that larger models perform comparably or even better than previously reported fine-tuned models on out-of-domain data. Second, we focus on Sanskrit by introducing a factoid closed-book QA dataset and show that question-answering performance improves with retrieval augmented generation (RAG) (Lewis et al., 2020) when the model is provided with relevant texts. Given the recent nature of these datasets relative to the models’ knowledge cut-off dates—and the likelihood that instruction-tuning on these languages is limited—the robust performance observed can be attributed to cross-lingual generalization. In both experimental setups, we find that smaller models struggle, particularly with niche entity types in NER, and in effectively leveraging contextual information via RAG suggesting model scale as an important factor.
@@ 2 Datasets and Methods
Task Lang. Source Size
@@
We evaluate the zero-shot capabilities of both large and small variants of our models: proprietary (gpt-4o and gpt-4o-mini (OpenAI, 2024)) and open-source (llama-3.1-405b-instruct and llama-3.1-8b-instruct (Dubey et al., 2024)). According to official documentation, these models have knowledge cut-off dates at the end of 2023. Given that none of the documentation indicates explicit instruction tuning on Sanskrit, Ancient Greek, or Latin, any observed performance in these languages is attributable to cross-lingual generalization. Previous zero-shot applications of LLMs to classical languages have been limited to Latin machine translation and summarization (Volk et al., 2024). All prompts used for these tasks are provided in Appendix C. The prompts are in English but also tested with manually translated san prompts for san. For the NER task, the prompt includes a single example to illustrate the output format, though it is effectively zero-shot for most entity types, as the example contains at most two.
@@ 2.1 Sanskrit QA
To further evaluate comprehension, we focus on question-answering (QA) in Sanskrit—a domain with very limited datasets. The only existing dataset by Terdalkar and Bhattacharya (2019) comprises 80 kinship-related questions. To address this gap, we created a new dataset containing 1501 factoid QA pairs covering distinct domains in Sanskrit: epics and healthcare. Specifically, we selected two key texts: (1) Rāmāyaṇa, an ancient Indian epic, and (2) Bhāvaprakāśanighaṇṭu, a foundational text on Āyurveda. Details of the dataset are provided in Appendix A.
For QA evaluation, we employ a closed-book setting using prompts detailed in Appendix C.3. Additionally, to simulate extractive QA, we implement Retrieval Augmented Generation (RAG) by retrieving the top-$k$ relevant passages (with $k$ tuned to 4) from the original Sanskrit texts using BM25 retrieval (Sparck Jones, 1972; Robertson et al., 2009). We found BM25 based RAG (RAG-BM25) to outperform RAG based on other possible retrievers that employ word-embeddings namely, FastText (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014) (reported in Appendix B). To determine if the presence of the answer in the context benefits the LLMs, we manually annotated the relevance of the retrieved passages. Since BM25 computes similarities between queries and texts preferably based on lemmas rather than inflected forms, a lemmatizer is needed in the pre-processing step. For this purpose, we trained a Seq2Seq transformer-based Sanskrit lemmatizer based on the DCS corpus (Hellwig, 2010-2024), achieving a mean F1-score of 0.94 on a held-out test set. Implementation details are given in Appendix D. The code and data are available at https://github.com/mahesh-ak/SktQA.
@@ 3 Results
Figure 1: Zero-shot evaluation of LLMs on three NLU tasks for classical languages (language codes in ISO 639-2). Prompts are in English except that <san> represents prompt in san. Larger LLMs are represented in red and orange, while smaller LLMs in blue and purple. First row displays the performances on NER (a) and MT (to en) (b) for all three languages. Second row displays the performances on QA for Sanskrit alone. Out of 1501 QA pairs whose chunks are considered while plotting (c), there are 607 QA pairs with answer present in at least one of the $k=4$ contexts extracted by BM25 and 894 QA pairs with answer not inferable from contexts, which are respectively considered in (d) and (e).
Figure 1 presents our zero-shot evaluation results, demonstrating that larger LLMs exhibit robust cross-lingual generalization across four NLU tasks—named entity recognition (NER), machine translation (MT), closed-book QA, and extractive QA (simulated via RAG-BM25)—in three classical languages (with QA evaluated solely on Sanskrit). To assess variability, each test set is segmented into 10 chunks during evaluation resulting in a box-plot (Fig. 1). Larger models perform comparable or better than previous fine-tuned models on out-of-domain data as reported in Appendix E. Notably, when answer-bearing contexts are provided (Fig. 1d) versus when they are absent (Fig. 1e), the models show significant performance gains, suggesting comprehension abilities in Sanskrit, a language characterized by high inflection. This behavior is however, not exhibited by smaller models when prompted in Sanskrit.
3.1 Sanskrit vs English Prompts
In our evaluations for san, we prompt in both English and Sanskrit, as mentioned earlier. Figure 1 shows that, for most models, prompting in English significantly outperforms prompting in Sanskrit—except for gpt-4o, where performance is largely similar. This provides partial evidence that these models are not instruction tuned on Sanskrit, aligning with trends observed by Muennighoff et al. (2023). In the case of gpt-4o, the comparable performance may indicate that the model has reached a saturation point to its best possible zero-shot performance as it is unlikely to have been trained on substantially more instruction tuning data than gpt-4o-mini.
Figure 2: Confusion matrices from the NER task in san (a-d), lat (e-h) and grc (i-l), normalized across rows.
@@ 3.2 Misclassified Entities in NER
Figure 2 displays confusion matrices for the NER task. Across all three languages, the smaller models exhibit more confusion among semantically related classes, while the larger models show fewer off-diagonal errors. In san, mythological entities like DEVA, ASURA, and RAKSHASA or semantically closed entities like KINGDOM vs CITY (eg., kośala vs ayodhyā) or FOREST (eg., daṇḍaka) vs GARDEN (eg., nandana, a heavenly garden)often get misclassified with each other in the smaller models. For lat, entity type GRP proves troublesome for the smaller models, suggesting struggles in separating individual entities from collective ones. In grc, categories such as LOC and ORG show higher confusion counts akin to GRP in lat while confusion between GOD and PERSON seems similar to what we have just discussed in Sanskrit. In contrast, much clearer boundaries emerge in the larger models’ confusion matrices. In many of these cases, the domain or style of the texts—especially if they involve mythological or archaic terms typical of classical texts— can be understood to influence performance, as models with less exposure to specialized terminology may conflate related entity types.
3.3 Inflection in Answers of san QA
@@
In the Sanskrit QA task, models are expected to generate single-word answers with correct inflection. For computing exact match (EM) scores, we manually identified all acceptable answers, excluding those with incorrect inflection (e.g., wrong case or gender endings). To quantify inflection errors, we also computed EM scores on lemmatized versions of both the gold-standard and predicted answers, as shown in Table 2. Most models exhibit only a slight increase in EM scores on lemmatized answers, suggesting that inflection errors are relatively minor—a finding corroborated by manual inspection. Future work could extend this analysis to investigate inflection accuracy in full sentence generation within broader Natural Language Generation scenarios.
3.4 Knowledge-Graph QA
Additionally, we explored the use of incomplete knowledge graphs (KGs) for Sanskrit QA. We evaluated a KG derived from the Bhāvaprakāśanighaṇṭu text (Terdalkar et al., 2023) and constructed a small KG for Rāmāyaṇa (details in Appendix F). Using the Think-On-Graph (ToG) paradigm (Sun et al., 2024), which iteratively explores KG paths for answer retrieval in a zero-shot, training-free manner (Xu et al., 2024), we observed that gpt-4o could effectively execute this method. Although it occasionally extracted correct answers, its performance did not significantly exceed that of the closed-book setting, possibly due to the incompleteness of the KGs. Future work may focus on developing more comprehensive KGs to enhance this retrieval method.
@@ 4 Conclusions
In summary, our zero-shot evaluations demonstrate that larger language models exhibit robust cross-lingual generalization across diverse natural language understanding tasks in classical languages—including NER, machine translation, and QA. Notably, the significant performance gains achieved when answer-bearing contexts are provided—particularly in Sanskrit QA—suggest comprehension abilities in highly inflected languages. Moreover, our contribution of a novel Sanskrit QA dataset provides a valuable resource for evaluating and benchmarking LLM performance on classical language tasks. Importantly, these models have not been explicitly instruction tuned on Sanskrit, Latin, or Ancient Greek—evidenced by the superior performance achieved when using English prompts for Sanskrit—which indicates that their zero-shot performance is attributable solely to cross-lingual generalization. Future work should focus on expanding dataset coverage and exploring additional classical languages and tasks, further advancing our understanding of cross-lingual generalization in LLMs and its applications in digital humanities and multilingual NLP research.
@@
While our study demonstrates robust cross-lingual generalization in large language models for classical languages, several limitations warrant discussion. First, our newly contributed Sanskrit QA dataset, although valuable, is limited in size. Our evaluation relies exclusively on zero-shot performance, as the models have not been explicitly instruction tuned on these languages; this design choice may obscure potential benefits achievable through targeted fine-tuning. Furthermore, the effectiveness of our BM25-based retrieval approach depends heavily on preprocessing steps such as lemmatization, which might not optimally address all linguistic variations in highly inflected languages. Finally, our comparisons are based on a limited set of proprietary and open-source models, and future work should extend this analysis to a broader range of models and tasks to fully understand the nuances of cross-lingual generalization in classical languages.
@@
Classical Sanskrit epics hold deep cultural and religious significance in Indian traditions, and similarly, Āyurveda represents a sensitive area within healthcare. We acknowledge that any research involving these subjects must be conducted with particular care. It is essential to note that, as with conventional treatment, Āyurvedic practices require professional consultation and should not be substituted by automated responses. Although our experiments indicate that paradigms like RAG produce more grounded and, hence, potentially safer outputs, there is no assurance that the responses from current LLMs in these domains meet clinical or religious safety standards. Accordingly, the authors do not endorse using the datasets beyond the scope of linguistic research. These datasets will be released under an open-source, non-commercial license, and all annotators have been compensated at fair, standard rates.
Appendix A Question Answering Dataset

### v2

@@ 1 Introduction
Large Language Models (LLMs) (Brown, 2020; Ouyang et al., 2022; Touvron et al., 2023) are known to generalize across various tasks using data from languages present in their pre-training phase, even when not present in instruction tuning (Wang et al., 2022; Muennighoff et al., 2023; Han et al., 2024). Previous work has demonstrated generalization to several low-resource languages via few-shot in-context learning (Cahyawijaya et al., 2024). In this study, we focus on zero-shot generalization to natural language understanding (NLU) tasks for three classical languages—Sanskrit (san), Ancient Greek (grc), and Latin (lat)—with a primary focus on Sanskrit. These languages represent a unique category of low-resource languages – despite scarce data for downstream NLU tasks, they have rich ancient literature available in digitized formats (Goyal et al., 2012; Berti, 2019), and they exert significant influence on the vocabulary and narrative structures of better-resourced languages (e.g., Latin contributes approximately 28% of English vocabulary (Finkenstaedt and Wolff, 1973)). Moreover, the high inflection of these languages presents a challenge.
Figure 1: The three NLU tasks evaluated on the classical languages: Named-Entity Recognition (top-left), Machine Translation (bottom-left) and Question-Answering (right).
To investigate these issues, we conduct two sets of zero-shot experiments using gpt-4o (OpenAI, 2024; OpenAI et al., 2023), llama-3.1-405b-instruct (Dubey et al., 2024), and their smaller variants. First, we assess performance on two NLU tasks with available datasets for all three languages, namely, named entity recognition (NER) and machine translation to English (MT). We observe that larger models perform comparably or even better than previously reported fine-tuned models on out-of-domain data. Second, we focus on Sanskrit by introducing a factoid closed-book QA dataset and show that question-answering performance improves with retrieval augmented generation (RAG) (Lewis et al., 2020) when the model is provided with relevant texts. The tasks are illustrated in Figure 1.
Given the recent nature of these datasets relative to the models’ knowledge cut-off dates, and the likelihood that instruction-tuning on these languages is limited, the robust performance observed can be attributed to cross-lingual generalization. We refer to our prompting strategy as zero-shot, as it includes no task-specific examples, and it is unlikely that such examples in these languages were present in the models’ training or instruction-tuning data. In both experimental setups, we find that smaller models struggle, particularly with niche entity types in NER, and in effectively leveraging contextual information via RAG, thereby suggesting model scale as an important factor.
@@ 2 Datasets and Methods
Task Language Source Size
@@
We evaluate the zero-shot capabilities of both large and small variants of our models: proprietary (gpt-4o and gpt-4o-mini (OpenAI, 2024)) and open-source (llama-3.1-405b-instruct and llama-3.1-8b-instruct (Dubey et al., 2024)). According to official documentation, these models have knowledge cut-off dates at the end of 2023. Many datasets considered in this work (Table 1) are released beyond these dates, in other words, they are unlikely to be seen in the pre-training data of these models. Given that none of the documentation indicates explicit instruction tuning on Sanskrit, Ancient Greek, or Latin, any observed performance in these languages is likely attributable to cross-lingual generalization. Previous zero-shot applications of LLMs to classical languages have been limited to Latin machine translation and summarization (Volk et al., 2024), although several works have been dedicated to building language models for these languages (Riemenschneider and Frank, 2023; Nehrdich et al., 2024), however, with fine-tuning restricted to morphological parsing-related tasks like dependency parsing (Nehrdich and Hellwig, 2022; Hellwig et al., 2023; Sandhan et al., 2023).
All prompts used for these tasks are provided in Appendix A. The prompts are tested in both English and the respective languages.
@@ 2.1 Sanskrit QA
To further evaluate comprehension, we focus on question-answering (QA) in Sanskrit – a domain with very limited datasets. The only existing dataset by Terdalkar and Bhattacharya (2019) comprises 80 kinship-related questions. To address this gap, we created a new dataset containing 1501 factoid QA pairs covering distinct domains in Sanskrit: epics and healthcare. Specifically, we selected two key texts: (1) Rāmāyaṇa, an ancient Indian epic, and (2) Bhāvaprakāśanighaṇṭu, a foundational text on Āyurveda. Details of the dataset are provided in Appendix B.
For QA evaluation, we employ a closed-book setting using prompts detailed in Appendix A.3. To emulate extractive QA, we implement a Retrieval-Augmented Generation (RAG) approach by retrieving the top-$k$ relevant passages ($k$ tuned to 4) from the original Sanskrit texts using BM25 (Sparck Jones, 1972; Robertson et al., 2009). We also compare BM25 with embedding-based retrievers—FastText (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014)—and vary $k$ to assess performance using gpt-4o with Sanskrit prompts. As shown in Fig. 2, BM25 consistently outperforms the embedding-based methods, and $k=4$ emerges as an optimal choice across metrics.
To examine whether the inclusion of answer-bearing contexts benefits model performance, we manually annotated the relevance of retrieved passages. Since BM25 relies on lexical similarity, typically favoring lemmas over inflected forms, we introduce a lemmatization step using a transformer-based Seq2Seq Sanskrit lemmatizer trained on the DCS corpus (Hellwig, 2010-2024), achieving a mean F1 score of 0.94 on a held-out test set. Further details on RAG and lemmatization are provided in Appendix C, and implementation details in Appendix D. Code and data are available at https://github.com/mahesh-ak/SktQA.
@@ 3 Results
Figure 3: Zero-shot evaluation of LLMs on three NLU tasks for classical languages (language codes in ISO 639-2). Prompts when in English are denoted by <en>, otherwise are in respective languages. Larger LLMs are represented in red and orange, while smaller LLMs in blue and purple. First row displays the performances on NER (a) and MT (to en) (b) for all three languages. Second row displays the performances on QA for Sanskrit alone. Out of 1501 QA pairs considered (c), 607 QA pairs are with answer present in at least one of the $k=4$ contexts extracted by BM25 and 894 QA pairs with answer not inferable from contexts, which are respectively considered in (d) and (e).
Figure 3 presents our zero-shot evaluation results, demonstrating that larger LLMs exhibit robust cross-lingual generalization across four NLU tasks—named entity recognition (NER), machine translation (MT), closed-book QA, and extractive QA (simulated via RAG-BM25)—in three classical languages (with QA evaluated solely on Sanskrit). To assess variability, each test set is segmented into 10 chunks during evaluation resulting in a box-plot. Larger models perform better than previous fine-tuned models on out-of-domain data as reported in Appendix E. Notably, when answer-bearing contexts are provided (Fig. 3d) versus when they are absent (Fig. 3e), the models show significant performance gains, suggesting comprehension abilities in Sanskrit, a language characterized by high inflection. This behavior is however, not exhibited by smaller models when prompted in Sanskrit.
3.1 Prompt Language: English versus Native
During evaluation, we prompted models both in English and in each target language. As shown in Figure 3, English prompts generally outperform Sanskrit prompts, particularly with smaller models, providing partial evidence that these models have not been instruction-tuned on Sanskrit (Muennighoff et al., 2023). For Latin and Ancient Greek, this English-prompt advantage holds mainly for smaller models; larger models perform equally well, or even better, with native-language prompts (e.g., in Latin NER). This does not imply instruction tuning in these languages, since larger and smaller models likely saw comparable amounts of tuning data. Rather, it suggests that cross-lingual transfer is especially strong for Latin and Ancient Greek in larger models, reflecting their substantial influence on high-resource languages such as English.
Figure 4: Confusion matrices from the NER task in san (a-d), lat (e-h) and grc (i-l), all with <en> prompts, normalized across rows.
@@ 3.2 Misclassified Entities in NER
Figure 4 displays confusion matrices for the NER task. Across all three languages, the smaller models exhibit more confusion among semantically related classes (see G for descriptions of entity types), while the larger models show fewer off-diagonal errors. In san, mythological entities like Deva, Asura, and Rakshasa or semantically closed entities like Kingdom versus City (e.g., Kośala vs Ayodhyā) or Forest (e.g., Daṇḍaka) versus Garden (e.g., Nandana) often get misclassified with each other in the smaller models. For lat, entity type GRP proves troublesome for the smaller models, suggesting struggles in separating individual entities from collective ones. In grc, categories such as LOC and ORG show higher confusion counts akin to GRP in lat while confusion between God and Persons seems similar to what was discussed for Sanskrit. In contrast, much clearer boundaries emerge in the larger models’ confusion matrices. In many of these cases, the domain or style of the texts, especially if they involve mythological or archaic terms typical of classical texts, can be understood to influence performance, as models with less exposure to specialized terminology may conflate related entity types.
3.3 Inflection in Answers in Sanskrit QA
@@
In the Sanskrit question-answering task, models are expected to generate single-word answers with the correct inflection. For computing exact match (EM) scores, we manually identified all acceptable answers, excluding those with incorrect inflection (e.g., wrong case or gender endings). To quantify inflection errors, we also calculated EM scores on lemmatized versions of the gold standard and predicted answers, as shown in Table 2. Most models show only a slight increase in EM scores on lemmatized answers, suggesting that inflection errors are relatively minor, a finding corroborated by manual inspection. Future work could extend this analysis to investigate inflection accuracy in full sentence generation within broader natural language generation scenarios.
3.4 Sanskrit Orthography: Devanagari versus IAST
LLM MT (BLEU) NER (Macro F1-BI)
Devanagari IAST Devanagari IAST
gpt-4o 0.179 0.165 0.637 0.599
llama-v3p1-405b-instruct 0.193 0.148 0.561 0.556
gpt-4o-mini 0.135 0.099 0.359 0.318
llama-v3p1-8b-instruct 0.120 0.063 0.164 0.149
Table 3: Comparison of performances in san MT and NER (English prompts) when the input sentences are given Devanagari script or in IAST script.
So far, we have shown robust cross-lingual generalization in the models. We now turn to one possible underlying mechanism—orthographic transfer—where models benefit from shared scripts across languages. Prior work has identified orthography as a key factor in cross-lingual transfer for LLMs (Muller et al., 2021; Fujinuma et al., 2022). To isolate this effect, we re-ran our Sanskrit NER and MT experiments (using English prompts) in Roman-based IAST transliteration instead of Devanagari. Table 3 compares performance in both scripts. Models perform better with the Devanagari script, which is shared by higher-resource relatives like Hindi and Marathi, reinforcing the importance of script sharing. However, results in IAST are only slightly lower, suggesting that Roman-based transliterations also feature prominently in the pre-training data. In future, we will investigate whether model outputs are consistent across both scripts, that is, whether these LLMs are effectively digraphic in Sanskrit.
3.5 Knowledge-Graph Question-Answering
Additionally, we explore the use of knowledge graphs (KGs) for Sanskrit QA. We evaluated a KG derived from the Bhāvaprakāśanighaṇṭu text (Terdalkar et al., 2023) and constructed a small KG for Rāmāyaṇa (details in Appendix F). Using the Think-On-Graph (ToG) paradigm (Sun et al., 2024), which iteratively explores the KG paths for answer retrieval in a training-free zero-shot manner (Xu et al., 2024), we observed that gpt-4o could effectively execute this method. Although it occasionally extracted correct answers, its performance did not significantly exceed that of the closed-book setting, most likely due to the incompleteness of the KGs (see §F.3). Future work may focus on developing more comprehensive KGs to enhance this retrieval method.
@@ 4 Conclusions
In summary, our zero-shot evaluations demonstrate that larger language models exhibit robust cross-lingual generalization across diverse natural language understanding tasks in classical languages, including NER, machine translation, and QA. Notably, the significant performance gains achieved when answer-bearing contexts are provided, particularly in Sanskrit QA, suggest comprehension abilities in highly inflected languages. Moreover, our contribution of a novel Sanskrit QA dataset provides a valuable resource for evaluating and benchmarking LLM performance on classical language tasks. Importantly, these models have not been explicitly instruction tuned on Sanskrit, Latin, or Ancient Greek—evidenced by the superior performance achieved when using English prompts for Sanskrit—which indicates that their zero-shot performance is attributable solely to cross-lingual generalization.
Future work will focus on expanding dataset coverage, knowledge graphs and exploring additional classical languages and tasks, further advancing our understanding of cross-lingual generalization in LLMs and its applications in digital humanities and multilingual NLP research.
Acknowledgements
This research is financially supported by the Indian Knowledge Systems (IKS) Division of Ministry of Education, Govt. of India (project number AICTE/IKS/RFP1/2021-22/12). Mahesh Akavarapu received funding from Volkswagen Foundation under the Phylomilia project within the Pioneering Projects funding line. We also thank anonymous reviewers and the Area Chairs for their comments that have helped improve the paper.
@@ Acknowledgements
While our study demonstrates robust cross-lingual generalization in large language models for classical languages, several limitations warrant discussion. First, our newly contributed Sanskrit QA dataset, although valuable, is limited in size. Our evaluation relies exclusively on zero-shot performance, as the models have not been explicitly instruction tuned on these languages; this design choice may obscure potential benefits achievable through targeted fine-tuning. Further, a few datasets we experimented were released within the models’ knowledge cut-off dates raising the issue of data contamination. Among these, only Ancient Greek MT exhibits anomalously high performance, suggesting possible exposure. In general, NER, owing to its structural data should be less susceptible to contamination than MT. Furthermore, the effectiveness of our BM25-based retrieval approach depends heavily on preprocessing steps such as lemmatization, which might not optimally address all linguistic variations in highly inflected languages. Finally, our comparisons are based on a limited set of proprietary and open-source models, and future work should extend this analysis to a broader range of models and tasks to fully understand the nuances of cross-lingual generalization in classical languages.
@@
Classical Sanskrit epics hold deep cultural and religious significance in Indian traditions, and similarly, Āyurveda represents a revered tradition-bound area within healthcare. We acknowledge that any research involving these subjects must be conducted with particular care. It is essential to note that, as with conventional treatment, Āyurvedic practices require professional consultation and should not be substituted by automated responses. Although our experiments indicate that paradigms like RAG produce more grounded and, hence, potentially safer outputs, there is no assurance that the responses from current LLMs in these domains meet clinical or religious safety standards. Consequently, the authors do not endorse using the datasets beyond the scope of linguistic research. These datasets are released for open-source, non-commercial use, and all annotators have been compensated at fair, standard rates.
Appendix
Appendix A Prompts
The Sanskrit prompts are in Devanagari script. In this appendix, we provide these prompts transliterated in IAST scheme.
A.1 Prompts for Named Entity Recognition
Prompt in English
Recognize the named entities from the following sentence in {LANGUAGE}. The valid tags are {ENTITY TYPES}. Do not provide explanation and do not list out entries of ‘O’. Example:
Sentence: <word_1> <word_2> <word_3> <word_4> <word_5>
Output: {{‘B-<entity1>’: [‘<word_1>’, ‘<word_4>’], ‘B-<entity2>’:[‘<word_5>’]}}
Sentence: {INPUT}
Output:
(The example is never a real sentence and is only provided to specify the output structure. Hence, the evaluations are strictly zero-shot.)
Prompt in Sanskrit
adho datta vākye nāmakṛtāḥ sattvāḥ (named entities) abhijānīhi. tadapi vivṛtam mā kuru, kevalam pṛṣṭa viṣayasya uttaram dehi. api ca ‘O’-sambandhitāni na deyāni.
sattvāḥ eteṣu vargeṣu vartante - {ENTITY TYPES}. udāharaṇāya -
vākyam: <padam_1> <padam_2> <padam_3> <padam_4> <padam_5>
phalitam: {{ ‘B-<sattvaḥ1>’: [‘<padam_1>’, ‘<padam_4>’], ‘B-<sattvaḥ2>’: [‘<padam_5>’]}}
vākyam: {INPUT}
phalitam:
Prompt in Latin
Agnosce nomina propria (named entities) ex hac sententia Latina. Notae validae sunt {ENTITY TYPES}. Explanationem ne praebeas nec elementa ‘O’ elenca. Exemplar:
Sententia: <verbum_1> <verbum_2> <verbum_3> <verbum_4> <verbum_5>
Productus: {{‘B-<entitatem1>’: [‘<verbum_1>’, ‘<verbum_4>’], ‘I-<entitatem1>’: [‘<verbum_2>’], ‘B-<entitatem3>’:[‘<verbum_5>’]}}
Sententia: {INPUT}
Productus:
Prompt in Ancient Greek
[…]
$\pi\rho\acute{o}\tau\alpha\sigma\iota\varsigma$: <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_1> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_2> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_3> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_4> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_5>
[…]
$\pi\rho\acute{o}\tau\alpha\sigma\iota\varsigma$: {INPUT}
$\pi\alpha\rho\alpha\gamma\omega\gamma\acute{\eta}$:
A.2 Prompts for Machine Translation
Prompt in English
Prompt in Sanskrit
Prompt in Latin
Verte hanc sententiam Latinam in Anglicam. Nullam explicationem praebe.
Prompt in Ancient Greek
[…]
(Sanskrit QA Prompts)
In the following prompts, TOPIC is either ‘Rāmāyaṇa’ or ‘Āyurveda’.
A.3 Prompts for Closed-book QA
Prompt in English
{QUESTION} {CHOICES}
Prompt in Sanskrit
{QUESTION} {CHOICES}
A.4 Prompts for RAG-QA
Prompt in English
contexts: {CONTEXTS}
question:{QUESTION} {CHOICES}
Prompt in Sanskrit
sandarbhāḥ: {CONTEXTS}
praśnaḥ: {QUESTION} {CHOICES}
Appendix B Question Answering Dataset